OpenStudy (anonymous):

Anyone know how to find the vertex of f(x)=2x^2+4x-16 ?

OpenStudy (whpalmer4):

when you have the equation in standard form (as you do) y = ax^2 + bx + c the x value of the vertex is at x = -b/(2a) the y value of the vertex you find by plugging that value of x into the equation

OpenStudy (whpalmer4):

we have a = 2, b = 4, c = -16 x = -(4)/(2*2) = -4/4 = -1 what will the y value be? y = 2(-1)^2 + 4(-1) - 16 =

OpenStudy (whpalmer4):

you can also find the vertex by rearranging into vertex form, which is y = a(x-h)^2 + k the vertex will be at (h,k)

OpenStudy (anonymous):

factor first. 2x^2 + 8x - 4x -16 2x (x+4) -4 (x + 4) y= (2x-4)(x+4) your 2 roots are 2x-4= 0 2x=4 x=2 and x+4 = 0 x=-4 therefore h = 2-4 /2 h= -1 sub -1 as your x values into the equation to find y y= (2(-1) -4) (-1 +4) y= (-6)(3) y=-18

OpenStudy (anonymous):

therefore your vertex is (-1, -18)